OpenStudy (anonymous):

Solve for x square root of 2x-3+16=19 show steps?

myininaya (myininaya):

\[\sqrt{2x-3}+16=19 ?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

Apply the following Steps: (1) Subtract 16 on both sides (2) To get rid the square root on the left hand side, square both sides (3) Add 3 on both sides (4) Divide both sides by 2

myininaya (myininaya):

\[\sqrt{2x-3}+16-16=19-16\] applied first step \[\sqrt{2x-3}=3\] \[(\sqrt{2x-3})^2=3^2\] applied second step \[2x-3=9\] now see if you can finish this

OpenStudy (anonymous):

the answer is 6

myininaya (myininaya):

yep thats right gj